# Enigmatic factor of 4/3 in electromagnetic momentum of a moving spherical capacitor

### v3

@@ II.1 Energy-momentum computed from a 4-vector Lorentz transformation from the rest frame
The electromagnetic momentum $\bf P^{\prime}$ is zero in rest frame ${\cal K}^{\prime}$ as the magnetic field is zero throughout in that frame. If the system has an acceleration ${\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$ in the instantaneous rest frame, then we can write ${\rm d}{\bf P}^{\prime}/{\rm d}t^{\prime}=(U^{\prime}/c){\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$. The rate of change of energy ${U}^{\prime}$ in the rest frame is nil, i.e., ${\rm d}U^{\prime}/{\rm d}t^{\prime}=0$.
@@
Because of an acceleration ${\rm d}\mbox{\boldmath$\beta$}/{\rm d}t$ of the system in the lab frame, we have
@@ III.2 Energy contribution from electromagnetic forces during work done against Lorentz contraction
Suppose in the lab frame ${\cal K}$, where the capacitor system is moving with velocity $\beta=v/c$ along $x$-axis, the capacitor gets a velocity increment $\Delta\beta$ due to an acceleration component along the $x$-axis, and a corresponding increment in the Lorentz factor $\Delta\gamma\approx\gamma^{3}\beta\Delta\beta$, in a small time interval $\Delta t$. Because of the increase in the Lorentz factor, the distance between $R_{1}$ and $R_{2}$ decreases in frame ${\cal K}$ by
@@
Let the rod, as a result, gain a small velocity increment $\Delta\beta$ along $x$-axis, in lab frame ${\cal K}$. Now even if the push given to two ends of the rod is simultaneous in ${\cal K}^{\prime}$, in frame ${\cal K}$, where the rod is moving with velocity $\beta$, the trailing end of the rod will get the velocity increment $\Delta\beta$ before the leading end by a tiny time interval $l^{\prime}\gamma\beta/c$, thereby giving rise to a differential motion and a consequential shortening of the distance between the two ends by $l^{\prime}\gamma\beta\Delta\beta$, the same as obtained in Eq. (26), due to change in Lorentz contraction of the rod length. Thus we see that a Lorentz contraction is a real contraction in length due to relative movement of the system parts along the direction of motion. If there are outward forces on the rod ends, energy will be spent against these outward forces during the moving closer of two ends of the rod by a distance $l^{\prime}\gamma\beta/c$, in the lab frame.

### v4

@@ II.1 Energy-momentum computed from a 4-vector Lorentz transformation from the rest frame
The electromagnetic momentum $\bf P^{\prime}$ is zero in rest frame ${\cal K}^{\prime}$ as the magnetic field is zero throughout in that frame. If the system has an acceleration $c{\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$ in the instantaneous rest frame, then we can write ${\rm d}{\bf P}^{\prime}/{\rm d}t^{\prime}=(U^{\prime}/c){\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$. The rate of change of energy ${U}^{\prime}$ in the rest frame is nil, i.e., ${\rm d}U^{\prime}/{\rm d}t^{\prime}=0$.
@@
Because of an acceleration $c{\rm d}\mbox{\boldmath$\beta$}/{\rm d}t$ of the system in the lab frame, we have
@@ III.2 Energy contribution from electromagnetic forces during work done against Lorentz contraction
Suppose in the lab frame ${\cal K}$, where the capacitor system is moving with velocity $v=\beta c$ along $x$-axis, the capacitor gets a velocity increment $c\Delta\beta$ due to an acceleration component along the $x$-axis, and a corresponding increment in the Lorentz factor $\Delta\gamma\approx\gamma^{3}\beta\Delta\beta$, in a small time interval $\Delta t$. Because of the increase in the Lorentz factor, the distance between $R_{1}$ and $R_{2}$ decreases in frame ${\cal K}$ by
@@
Let the rod, as a result, gain a small velocity increment $c\Delta\beta$ along $x$-axis, in lab frame ${\cal K}$. Now even if the push given to two ends of the rod is simultaneous in ${\cal K}^{\prime}$, in frame ${\cal K}$, where the rod is moving with velocity $c\beta$, the trailing end of the rod will get the velocity increment $c\Delta\beta$ before the leading end by a tiny time interval $l^{\prime}\gamma\beta/c$, thereby giving rise to a differential motion and a consequential shortening of the distance between the two ends by $l^{\prime}\gamma\beta\Delta\beta$, the same as obtained in Eq. (26), due to change in Lorentz contraction of the rod length. Thus we see that a Lorentz contraction is a real contraction in length due to relative movement of the system parts along the direction of motion. If there are outward forces on the rod ends, energy will be spent against these outward forces during the moving closer of two ends of the rod by a distance $l^{\prime}\gamma\beta\Delta\beta$, in the lab frame.